TOWARD A CONTINUOUS-WAVE SOLID HYDROGEN RAMAN LASER FOR MOLECULAR SPEC-TROSCOPY APPLICATIONS

\underline{W. R. EVANS}, Department of Physics, University of Illinois at Urbana-Champaign, Urbana, IL 61801; T. MOMOSE, Department of Chemistry, The University of British Columbia, Vancouver, BC Canada V6T 1Z1; B. J. McCALL, Departments of Chemistry, Physics, and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

We will present our recent work toward the construction of a continuous-wave solid para-H_{2} Raman laser for operation first in the visible and later in the mid-infrared. Solid para-H_{2} promises to be a good choice for the gain medium in a Raman laser due to its exceptionally high Raman gain coefficient. This not only presents a novel use of an interesting molecular system, but it also offers the potential for the first widely tunable laser source for high resolution spectroscopy in the 5-10 μm range. High resolution spectroscopy requires a tunable continuous-wave laser source. However, up until now, most work in using para-H_{2} as a Raman laser gain medium has taken place either with high power pulsed lasers or continuous-wave lasers which require ultra-high finesse cavities. We seek to take advantage of solid para-H_{2}’s high Raman gain coefficient to construct a continuous-wave Raman laser with a much lower finesse cavity (F ≈ 150). In this presentation, we will talk about our recent work in measuring the index of refraction of solid para-H_{2} in the wavelength range 430-1100 nm in preparation for building such a laser. Some details regarding the design and planning for this laser will also be discussed. Finally, current progress and anticipated work on the development of a continuous-wave solid para-H_{2} Raman laser will be presented.